OpenStudy (anonymous):

Which is a counterexample to the conjecture? Any number that is divisible by 3 is divisible by 9. A. 18 B. 63 C. 27 D. 69

OpenStudy (anonymous):

@aloang33

OpenStudy (malcolmmcswain):

Wrong section.

OpenStudy (malcolmmcswain):

You're in Computer Science. But I'll still help.

OpenStudy (anonymous):

o sorry but thx you for helping me

OpenStudy (malcolmmcswain):

Which number is divisible by 3, but not by 9?

OpenStudy (anonymous):

idk

OpenStudy (malcolmmcswain):

Well 18/9 = ?

OpenStudy (anonymous):

2

OpenStudy (malcolmmcswain):

63/9 = ?

OpenStudy (anonymous):

7

OpenStudy (malcolmmcswain):

27/9 = ?

OpenStudy (anonymous):

3

OpenStudy (malcolmmcswain):

69/9 = ?

OpenStudy (anonymous):

621

OpenStudy (malcolmmcswain):

What? How did you get that?